ABC Costing for a Service Organization
Fairfield Mortgage Company is a full-service residential mortgage company in the Atlanta area that operates in a very competitive market. The CEO, Richard Sissom, is concerned about operating costs associated with processing mortgage applications and has decided to install an ABC costing system to help him get a handle on costs. Although labor hours seem to be the primary driver of the cost of processing a new mortgage, the labor cost for the different activities involved in processing new loans varies widely. The Accounting Department has provided the following data for the company's five major cost pools for 2019:

Taking customer applications $300,000

Time-assistant managers

12,000 hours
Conducting credit investigations 450,000

Time-credit managers

16,500 hours
Underwriting 525,000 Time-Underwriting Department 10,000 hours
Preparing loan packages 200,000

Time-Processing Department

8,000 hours
Closing loans 600,000

Time-Legal Department hours

6,000 hours
$2,075,000 52,500 hours

During 2019, the company processed and issued 5,000 new mortgages, two of which are summarized here with regard to activities used to process the mortgages:

Loan 5066 Loan 5429
Application processing hours 1.50 2.75
Credit investigating hours 4.00 3.00
Underwriting hours 2.50 4.75
Processing hours 3.50 3.00
Legal processing hours 1.50 1.50
Total hours 13.00 15.00

a. Determine the cost per unit of activity for each activity cost pool.
Round rates to two decimal places.
Activity-based overhead rates:

Taking applications Answer
Conducting credit investigations Answer
Underwriting Answer
Preparing loan packages Answer
Closing loans Answer

b. Determine the cost of processing loans 5066 and 5429.
Round answers to two decimal places.

Loan 5066 Loan 5429
Taking applications Answer Answer
Investigations Answer Answer
Underwriting Answer Answer
Loan packages Answer Answer
Closing loans Answer Answer
Total costs Answer Answer

c. Determine the cost of preparing loans 5066 and 5429 assuming that an average cost per hour for all activities is used.
Round answers to two decimal places.
Average cost per hour for all activities $Answer
Use rounded rate calculated above for calculations below. Round answers to two decimal places.

Cost of Loan #5066 $Answer
Cost of Loan #5429 $Answer
Answer #1

A)Cost per unit of activity for each activity cost pool:

ACTIVITY (1) ACTIVITY COST POOLS($) (2) COST DRIVERS (3) COST DRIVER QUANTITY(Hours) (4) COST PER UNIT OF ACTIVITY ($) {2 / 4}
Taking customer application 300,000 Application processing hours 12,000 25.00
Conducting credit investigation 450,000 Credit investigating hours 16,500 27.27
Underwriting 525,000 Underwriting hours 10,000 52.50
Preparing loan packages 200,000 Processing hours 8,000 25.00
Closing loans 600,000 Legal processing hours 6,000 100.00

B) Cost of processing loan 5066 & 5429:

PARTICULARS LOAN 5066 ($) LOAN 5429 ($)
Taking application

37.50

(25*1.5)

68.75

(25*2.75)

Incestigation

109.08

(27.27*4)

81.81

(27.27*3)

Underwriting

131.25

(52.50*2.5)

249.38

(52.50*4.75)

Loan packages

87.50

(25*3.5)

75.00

(25*3)

Closing loans

150.00

(100*1.50)

150.00

(100*1.50)

Total cost of processing 515.33 624.94

C)Average cost per hour for all activities:

It can be calculated as follows = Total(overall) processing costs Total number of hours for processing

= $2,075,000 / 52,500

= $39.52 per hour

Cost of loan 5066 = No. of hours*average cost per hour = 13*39.52 = $513.76

Cost of loan 5429 = No. of hours*average cost per hour = 15*39.52 = $592.80